Circular No. 7023

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

NOVAE IN M31 M. Papenkova, J. Y. King, E. Halderson, T. Shefler, M. Modjaz, W. D.

Li, R. R. Treffers, and A. V. Filippenko, University of California at Berkeley, report their discovery of an apparent nova in M31 (NGC 224) during the course of the Lick Observatory Supernova Search with the 0.8-m Katzman Automatic Imaging Telescope (KAIT). The nova was found and confirmed on unfiltered CCD images taken on Sept. 30.3 (mag ∼ 16.1) and Oct. 1.3 UT (mag ∼ 16.2). The object is located at α = 0^{h}42^{m}49^{s}.9, δ = +41^{o}16′48′′ (equinox 2000.0; due to a lack of nearby stars, this semiaccurate position was calculated as an offset from the NED position for the nucleus of M31), which is ∼ 63′′ east and 39′′ north of the galaxy nucleus. A KAIT image of M31 taken on July 23.5 shows nothing at the position of the nova (limiting mag ∼ 19.0). The nova reported on IAUC 7015 is also present on Sept. 30.3 and Oct. 1.3 KAIT frames at unfiltered mag ∼ 18.5.

SGR 1900+14 AND PSR J1907+09 K. Xilouris, National Radio Astronomy and Ionospheric Center; C.

Kouveliotou, Universities Space Research Association at Marshall Space Flight Center, NASA; D. R. Lorimer, Max Planck Institute for Radio Astronomy, Bonn; R. Ramachandran, University of Amsterdam (UoA); and J. van Paradijs, UoA and University of Alabama in Huntsville, report: “We observed the soft-γ-ray repeater SGR 1900+14 during June with the Arecibo telescope. A search for pulsed radio emission at 1.4 GHz did not detect the 5.16-s period reported for SGR 1900+14 (IAUC 7001). We have discovered, however, a 226-ms radio pulsar, PSR J1907+09. The nominal position of the pulsar is α = 19^{h}07^{m}21^{s}.1, δ = +9^{o}18′41′′ (equinox 2000.0), with an uncertainty of ∼ 2′. This is the first pulsar to be discovered with the recently upgraded Arecibo telescope. Confirmation observations at 1.4 GHz were made with the Effelsberg 100-m telescope several times during September. Measurements of the pulse period over a 112-day baseline give a period derivative of (1 ± 2) × 10^{−13} s/s (3σ upper limit). The initial estimate for the dispersion measure of 350 ± 10 pc cm^{−3} places the pulsar at a distance of ∼ 7000 pc. The mean flux density at 1.4 GHz is ∼ 0.3 mJy. Despite a number of observations with the Arecibo telescope at 430 MHz, the pulsar was so far not detected at this frequency, suggesting that the degree of interstellar scattering along this line-of-sight is large. We encourage further observations of the source.”

1998 October 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green